# Autodeleveraging: Impossibilities and Optimization

### v1

@@ Abstract
Autodeleveraging (ADL) is a last-resort loss socialization mechanism for perpetual futures venues. It is triggered when solvency-preserving liquidations fail. Despite the dominance of perpetual futures in the crypto derivatives market, with over \$60 trillion of volume in 2024, there has been no formal study of ADL. In this paper, we provide the first rigorous model of ADL. We prove that ADL mechanisms face a fundamental trilemma: no policy can simultaneously satisfy exchange solvency, revenue, and fairness to traders. This impossibility theorem implies that as participation scales, a novel form of moral hazard grows asymptotically, rendering ‘zero-loss’ socialization impossible. On the positive side, we show that three classes of ADL mechanisms can optimally navigate this trilemma to provide fairness, robustness to price shocks, and maximal exchange revenue. We analyze these mechanisms on the Hyperliquid dataset from October 10, 2025, when ADL was used repeatedly to close \$2.1 billion of positions in 12 minutes. By comparing our ADL mechanisms to the one used in practice on Hyperliquid and Binance, we demonstrate empirically that Hyperliquid’s production queue overutilized ADL by $\approx 8\times$ relative to our optimal policy, imposing roughly \$630 million of unnecessary haircuts on winning traders. This comparison also suggests that Binance overutilized ADL far more than Hyperliquid. Our results both theoretically and empirically demonstrate that optimized ADL mechanisms can dramatically reduce losses of trader profitability while maintaining exchange solvency.
@@ 1 Introduction
Perpetual futures (or simply, perpetuals) are by far the most popular crypto derivatives contract. These contracts allow for duration-independent hedging of cryptocurrencies, such as Bitcoin, Ethereum or Solana. That is, unlike traditional future contracts, perpetuals are not subject to a fixed expiration date. Instead, perpetuals are similar to contracts for difference, where market participants repeatedly send each other payments depending on how the spot price trades relative to the futures price. For instance, if the spot price is higher than the futures price, then traders with short futures positions pay the traders with long futures until the prices are equalized. We note that while Robert Shiller’s 1993 proposal for perpetual futures provides the genesis of this idea, the BitMEX exchange’s 2016 launch of the XBTUSD perpetual swap became the first live implementation [shiller1993measuring, Soska2021BitMEX, Hayes2025AdaptOrDie].
Perpetual futures have a higher notional trading volume than spot cryptocurrency volume. This is akin to index futures for US stocks, where there is far more volume traded in the futures contracts than the underlying spot assets, with the S&P 500 futures index future trading roughly ten times more volume than the spot asset [CMEGroup2024EqualWeight]. In 2024, centralized crypto exchanges such as Binance and Bybit processed nearly \$58.5 trillion (notional) in perpetuals trades [CoinGecko2025Perps]. On the other hand, centralized exchanges processed nearly \$17.4 trillion of spot cryptocurrency trades [CoinGecko2025Q1], which is a 3.3x ratio of perpetuals to spot volume.
The main reason for the elevated usage of perpetual futures, much like their index futures cousins, is due to the high level of leverage they offer. In particular, the capital requirements for trading perpetual futures are much lower than spot trading. Moreover, the continuous nature of perpetual futures allows for easier margin management than duration-based futures. This is illustrated via the large gap between the maximum leverage offered by perpetuals exchanges (up to 125x [Binance2025CollateralLeverageUpdate]) versus the index futures on the Chicago Mercantile Exchange (around 10-15x leverage for index futures [CMEGroup2025SP500Margins]).
@@
Historically, the majority of perpetuals trading volume has been concentrated on centralized exchanges (CEXes). These venues, which include Binance, Bybit, and Coinbase, require users to deposit collateral that is custodied by the exchange. The exchange is responsible for ensuring that users’ positions are solvent and that their collateral is not used to cover losses from other traders. In practice, this involves traders having to trust the exchange to not use their collateral for other purposes. In 2022, centralized exchange FTX was found to have utilized customer collateral for other purposes, leading to multi-billion-dollar losses [FTXDebtors2023Report].
In response, there has been an increasing trend in the usage of decentralized exchanges (DEXes) for trading perpetual futures. These exchanges provide far more transparency and auditability into the mechanics of the exchange. In particular, user’s can always verify what their collateral is being used for and what positions it is collateralizing. Moreover, these exchanges are permissionless and pseudonymous, meaning anyone can trade on these exchanges without explicitly revealing their identity.
@@
However, in recent years, hybrid models of decentralized exchanges where withdrawals and deposits are permissionless, but certain exchange operations are centralized, have become popular. The most popular exchange of this form is Hyperliquid, which has helped increase the market share of perpetuals volume on DEXs from roughly 1% in 2024 to over 8% in 2025. These hybrid models allow for the exchange to use cryptographic commitments to enforce collateral rights while also allowing for oracles, markets, and solvency to be controlled by a smaller set of participants. This allows the exchange to avoid issues that fully open decentralized exchanges have, where weak collateral can be added to an exchange that cannot be sold or liquidated successfully during a market crash to ensure solvency.
More concretely, there have been recent incidents on perpetual DEXs that highlight how novel market structures can be exploited or stressed in practice. These include but are not limited to: Hyperliquid’s XPL pre-market price spike and the JellyJelly attack [CryptoTimesXPL2025, ODailyXPL2025, OakResearchJELLY]. These episodes led to emergency safeguards and community post-mortems, underscoring the need for robust oracle design, listing policies, and circuit breakers to mitigate manipulation. Such emergency responses would have been nearly impossible in the fully decentralized setting, demonstrating the value of hybrid models of decentralization in perpetual futures.
@@
Despite the empirical prominence of perpetuals markets, there has been little formal study of the stability and robustness of these markets relative to the large body of research into spot cryptocurrency trading. For instance, the main price stability mechanism used for perpetuals is the funding rate, a continuous payment stream between the long and short positions. There has been some study of how to replicate funding rates via other financial instruments [AngerisChitra2023PerpsSIAM, he2022fundamentals, ackerer2024perpetual], but little to no empirical study of how well such approximations work. Even more surprising is the lack of formal study of how robust perpetual exchanges are under high leverage conditions, despite their usage being driven by high leverage traders.
A crucial component to ensuring the stability of perpetual futures markets with high leverage is liquidation. This is the process of closing a user’s position that is worth less than the cash collateral they have posted to the exchange. Perpetuals exchanges rely on liquidation mechanisms to ensure that they stay solvent — that is, the assets they hold are greater than the liabilities they owe to traders. To the author’s knowledge, there has been no formal study of liquidations in perpetuals exchanges, despite there being a large body of work studying liquidations in overcollateralized cryptocurrency lending (e.g. [perez2021liquidations, kao2020analysis, sun2022liquidity, qin2021empirical]).
@@
Autodeleveraging is unique in that it algorithmically socializes losses on winners based on a public rule or mechanism that an exchange posts. This means that all of the position closures and implicit liquidations of insolvent positions are performed in an irreversible and atomic manner. The closest analogue to autodeleveraging in traditional finance is a central counterparty (CCP) such as a derivatives clearing house. In traditional finance, CCPs employ default waterfalls and loss mutualization (including variation-margin gains haircutting (VMGH)) that share the same core idea of allocating residual losses across non-defaulting members [DuffieZhu2011, Pirrong2011, CPMI_IOSCO_2014, ContGhamamiSITG, KubitzaWinnersLosers, Turing2019MagicRelighting]. CCPs generally have manual, non-algorithmic socialization and disbursement of insurance funds during such events with disputes reconciled via the legal system. We note, furthermore, that only clearing members bare the losses in CCPs, which leads to a different principal-agent dynamic than perpetuals markets as smaller traders need not bare the final losses.
Autodeleveraging, to the best of the author’s knowledge, was first introduced by the crypto exchange Huobi in 2015 [HuobiADL]. Prior to the implementation of ADL, exchanges would manually socialize losses, with the exchange serving as the sole CCP. By 2016, the BitMEX exchange implemented ADL via a formula that ranked trader positions for ADL by their profit times their leverage [BitMEXADL]. This formula persists to today on virtually every centralized exchange (including Binance, who implemented the formula in 2019 [BinanceADL]) and on the decentralized exchange Hyperliquid [HyperliquidDocsLiquidations]. We note that there are some decentralized exchanges such as Drift [DriftADLCode] and Paradex [ParadexADLBlog] that use a different ADL mechanism (the pro-rata mechanism defined in §2.4). However, well over 95% of perpetuals trading volume is on exchanges that utilize the BitMEX ranking model.
@@
Inherently, ADL creates moral hazard — that is, losing traders have an incentive to take on more risk in anticipation of being socialized by winning traders. The goal of an exchange is to balance the risk of insolvency against haircutting winning trader profits. Moral hazard is measured by this trade-off — if it is possible to reduce the risk of insolvency to zero without haircutting winning traders, then the system has no moral hazard. On the other hand, if it is not possible to reduce the risk of insolvency by any appreciable amount without severely haircutting winning trades, the system has excessive moral hazard. The job of an exchange operator is to choose an ADL mechanism that tries to induce a low level of moral hazard under most tail event scenarios. This connects to principal–agent models of moral hazard and linear/robust contracting [Holmstrom1982, Carroll2015RobustLinearContracts, DuttingEtAl2023MultiAgentContracts], with key differences: many agents act simultaneously, externalities operate through a common solvency constraint, and tail-risk mitigation substitutes for standard effort observability.
@@
Despite the rarity of ADL, it is certainly not hypothetical: there have been dozens of instances of ADL activations on major venues since 2018 [BinanceADL, BitMEXADL, AevoADL, CoinglassADL]. The largest episode of ADL usage was on October 10-11, 2025 (UTC), when outsized liquidations exhausted insurance funds and multiple venues simultaneously invoked ADL [CoinDesk2025LargestLiquidations]. However, other famous incidents such as the March 2020 “Black Thursday” and May 19, 2021 deleveraging event were other prominent days involving ADL usage. Most of these usages of ADL often involved controversies, in that highly profitable but lower risk (e.g. lower leverage) traders found their positions closed out by ADL mechanisms. These traders have often threatened to sue exchanges for unfair treatment, with rumors and public speculation surfacing during Oct. 10–11, 2025 (subsequently denied by Wintermute) [CointelegraphWintermute2025, TheBlockWintermute2025] [^1] See also prior complaints during May 2020 on Binance futures when profitable shorts were closed via ADL [CointelegraphBinanceADL2020].
@@
This incidents demonstrate a natural trade-off that perpetuals exchanges have to make: they can either aggressively socialize losses to their winners and (potentially) lose the future revenue of these users while preserving solvency or hold the losses due to the insolvency on their balance sheet. In the former case, the exchange effectively chooses to prioritize short term solvency and potential liquidation profits over long term value and fees generated by users. The latter case, effectively creates a long term risk for the exchange’s balance sheet, especially if it does not have sufficient funds to cover insolvencies. This strongly suggests that one can view the choice of how an exchange implements ADL as a trade-off between solvency, moral hazard, and long term revenue for an exchange.
@@
We note that open interest post October 10, 2025 that suggests that Hyperliquid has nearly 50% of open interest, whereas competitors such as Lighter and even Binance have recovered to pre-event levels. [^2] Source: DeFiLlama perpetuals dashboard, https://defillama.com/perps, accessed Nov. 30, 2025. Open interest is generally more expensive to manipulate than trading volume and can serve as a coarse measurement of expected future revenue for an exchange. Numerous parties have argued that this open interest loss for Hyperliquid is due to their overly aggressive ADL policy. [^3] See, e.g. , public commentary from @fiddybps1, @0xReaz, @Eugene_Bulltime, and @0xLouisT on X (Nov. 2025): https://x.com/fiddybps1/status/1978750722901762321, https://x.com/0xReaz/status/1986486213101166599, https://x.com/Eugene_Bulltime/status/1994380900067582182, https://x.com/0xLouisT/status/1990815109787058654.
@@
    Empirical: Using Hyperliquid’s data from October 10, 2025, where ADL was used over 40 times in a 10 minute period, we evaluate the performance of difference ADL mechanisms
@@
Our negative results focus on quantifying limits to how much ADL can actually reduce insolvency without inuring traders with disproportionate losses. We explicitly ask the question of how the loss of profit due to ADL to the trader with the highest PNL scales as the number of positions $n$ grows. We quantify this by looking at ratios of the maximum winner’s profit to the total shortfall (over all traders) and to the worst trader’s shortfall. These ratios can be viewed as analogues of the Value-at-Risk and Expected Shortfall metrics in traditional risk modeling, as we show in Appendix B. The main difference is that while Value-at-Risk and Expected Shortfall measure the total expected loss (over all traders) in tail events, our metrics only measure the expected loss to the most winning trader. Moreover, Appendix B.5 shows that classical VaR/ES rankings echo the same conclusion: queue-based ADL always delivers the lowest top-winner survival even under these traditional risk lenses.
@@
As the negative results show that no ADL mechanism can scale to preserve winner’s profit, a natural follow-up question is, what is the most fair way of socializing profitable users? We consider two notions of fairness in §6: axiomatic and optimization-based. They are both common in the algorithmic fairness literature within machine learning [KleinbergEtAl2018FairnessImpossibility, BarocasHardtNarayanan2019FairnessBook, Mehrabi2021SurveyFairness, DworkEtAl2012FairnessAwareness, Agarwal2018ReductionsFairness, HardtPriceSrebro2016EqualityOfOpportunity]. We demonstrate that the pro-rata ADL rule (used by Drift [DriftADLCode] and Paradex [ParadexADLBlog]) is the unique fair ADL rule in both the axiomatic and optimization senses [johnson2023concave].
Axiomatic fairness involves defining rules or invariants that a particular ADL mechanism must satisfy to be deemed fair. The rules we analyze for fairness are monotonicity, scale invariance and Sybil resistance. Roughly speaking, monotonicity says that if you were the $k$th highest winner (ranked by PNL) before ADL, then you will be the $k$th highest winner after ADL. Scale invariance states that if everyone scaled up their positions by an equal factor, their loss due to ADL does not change. And finally, Sybil resistance states that splitting a position into multiple positions does not change the loss due to ADL. We show that under mild smoothness conditions, the pro-rata ADL rule is the unique rule to satisfy these axioms. In stark contrast, we show that the queue-based mechanisms used by Binance and Hyperliquid are effectively “anti-fair”: in our model and data they produce extremely low Profit-to-Solvency Ratios (PTSR), imposing disproportionately large losses on the top winner. This finding contradicts the common intuition that queues are orderly; instead, they concentrate losses on the most successful traders.
@@
A direct extension of the robustness threat model is to consider a sequential ADL problem, where for multiple rounds, a price shock is received, the exchange chooses an ADL policy, and then losses are realized. This model might at first glance seem to only have theoretical value since ADL is itself a rare event. However, on October 10, 2025 there were multiple ADL events that occurred over a short time window on both Hyperliquid and Binance. While the account-level data for Binance’s usage of ADL is not public, we do have this data for Hyperliquid. Starting at 10:21:00 UTC, there were at least 40 recorded ADL events on Hyperliquid in a 10 minute period. This strongly suggests that there is clustering or cascading of ADL events, making the multiple round model necessary to study.
@@
We first show that viewing ADL as a dynamic, multiple round Stackelberg game leads to sharply different equilibria. We prove a crucial “principal-agent” tension: any strategy an exchange executes that minimizes the time to recover solvency (i.e., number of rounds) necessarily maximizes the loss of exchange long-term revenue. This trade-off is the economic heart of the paper: the exchange wants to be solvent in few rounds, but doing so destroys the best customers (e.g. the whales). Intuitively, this occurs because any strategy that quickly recovers solvency autodeleverages the largest winners more aggressively. These winners, however, are also often the most profitable users and generate high long-term fees for the exchange. By liquidating them aggressively, you’re removing the capital that they would use to later create positions and generate fees. We note that threats to sue exchanges for ADL losses in profit after October 10, 2025 provide indirect empirical evidence to support this theoretical result [CointelegraphWintermute2025, TheBlockWintermute2025].
@@
This implies that any optimal ADL policy over multiple ronds also has to include incentive compatibility constraints that balance future exchange revenue, current exchange solvency, and current trader profits. We construct an explicit algorithm for optimizing a revenue and solvency based objective and show that a simple mirror descent algorithm achieves vanishing regret for this objective. We conclude with a simple example demonstrating that no static policy that dominate the dynamic policy, especially when exchange long-term revenue is a consideration.
Empirical Results: The \$630 Million Inefficiency.
We conclude by analyzing the October 10, 2025 ADL events on Hyperliquid. Unlike centralized exchanges like Binance and Bybit, Hyperliquid publicly shares all position data, including position leverage, collateral, PNL, and ADL status. Using this data, we evaluate the efficiency of Hyperliquid’s ADL mechanism (which is a slight modification of the original 2015 Huobi queueing mechanism). We show that Hyperliquid autodeleveraged roughly  \$705m of winning trader PNL during the 12 minute cascade on October 10, 2025. Over the same period, there were at least 40 ADL events which accumulated a total of  \$304m of negative equity positions. [^4] Hyperliquid’s on-chain fill log for 10 Oct. 2025 records 34,983 individual ADL executions across 19,337 distinct wallets and 162 tickers. We analyze the per-fill data but aggregate those executions into the 40 queue waves discussed in the text. Measured shock-by-shock, this implies the production queue generated about  \$630m of overshoot—haircuts beyond the deficits that actually materialized. This result highlights a massive capital inefficiency in current queue-based mechanisms: they are not just unfair; they destroy hundreds of millions of dollars of trader value unnecessarily.
We then compare the pro-rata, risk-weighted pro-rata, and dynamic mechanisms to the live queueing mechanism used. We utilize these methods with different severity optimization algorithms and show that all of these algorithms were able to learn to autodeleverage substantially less than what Hyperliquid did in practice. Utilizing these algorithms on the historical data, we find that nearly 98% of the over autodeleveraging used by Hyperliquid could have been removed by a smarter ADL policy.

### v3

@@ Abstract
Autodeleveraging (ADL) is a last-resort loss socialization mechanism for perpetual futures venues. It is triggered when solvency-preserving liquidations fail. Despite the dominance of perpetual futures in the crypto derivatives market, with over \$60 trillion of volume in 2024, there has been no formal study of ADL. In this paper, we provide the first rigorous model of ADL. We prove that ADL mechanisms face a fundamental trilemma: no policy can simultaneously satisfy exchange solvency, revenue, and fairness to traders. This impossibility theorem implies that as participation scales, a novel form of moral hazard grows asymptotically, rendering ‘zero-loss’ socialization impossible. On the positive side, we show that three classes of ADL mechanisms can optimally navigate this trilemma to provide fairness, robustness to price shocks, and maximal exchange revenue. We analyze these mechanisms on the Hyperliquid dataset from October 10, 2025, when ADL was used repeatedly to close \$2.1 billion of positions in 12 minutes. By comparing production ADL to transparent benchmark allocations, we find that Hyperliquid’s production algorithm overshot the minimum trader profit haircut required to cover the shortfall. Our methodology suggests the excess profits lost by profitable traders is between \$45.0M and \$51.7M. In terms of the positions liquidated, this corresponds to roughly \$653.6M of positions being closed. This comparison also suggests that Binance overutilized ADL far more than Hyperliquid. Our results show both theoretically and empirically that optimized ADL mechanisms can dramatically reduce losses of trader profitability while maintaining exchange solvency.
@@ 1 Introduction
Perpetual futures (or simply, perpetuals) are by far the most popular crypto derivatives contract. These contracts allow for duration-independent hedging of cryptocurrencies, such as Bitcoin, Ethereum or Solana. That is, unlike traditional future contracts, perpetuals are not subject to a fixed expiration date. Instead, perpetuals are similar to contracts for difference, where market participants repeatedly send each other payments depending on how the spot price trades relative to the futures price. For instance, if the spot price is higher than the futures price, then traders with short futures positions pay the traders with long futures until the prices are equalized. We note that while Robert Shiller’s 1993 proposal for perpetual futures provides the genesis of this idea, the BitMEX exchange’s 2016 launch of the XBTUSD perpetual swap became the first live implementation [106, 107, 65].
Perpetual futures have a higher notional trading volume than spot cryptocurrency volume. This is akin to index futures for US stocks, where there is far more volume traded in the futures contracts than the underlying spot assets, with the S&P 500 futures index future trading roughly ten times more volume than the spot asset [35]. In 2024, centralized crypto exchanges such as Binance and Bybit processed nearly \$58.5 trillion (notional) in perpetuals trades [38]. On the other hand, centralized exchanges processed nearly \$17.4 trillion of spot cryptocurrency trades [37], which is a 3.3x ratio of perpetuals to spot volume.
The main reason for the elevated usage of perpetual futures, much like their index futures cousins, is due to the high level of leverage they offer. In particular, the capital requirements for trading perpetual futures are much lower than spot trading. Moreover, the continuous nature of perpetual futures allows for easier margin management than duration-based futures. This is illustrated via the large gap between the maximum leverage offered by perpetuals exchanges (up to 125x [23]) versus the index futures on the Chicago Mercantile Exchange (around 10-15x leverage for index futures [36]).
@@
Historically, the majority of perpetuals trading volume has been concentrated on centralized exchanges (CEXes). These venues, which include Binance, Bybit, and Coinbase, require users to deposit collateral that is custodied by the exchange. The exchange is responsible for ensuring that users’ positions are solvent and that their collateral is not used to cover losses from other traders. In practice, this involves traders having to trust the exchange to not use their collateral for other purposes. In 2022, centralized exchange FTX was found to have utilized customer collateral for other purposes, leading to multi-billion-dollar losses [56].
In response, there has been an increasing trend in the usage of decentralized exchanges (DEXes) for trading perpetual futures. These exchanges provide far more transparency and auditability into the mechanics of the exchange. In particular, users can always verify what their collateral is being used for and what positions it is collateralizing. Moreover, these exchanges are permissionless and pseudonymous, meaning anyone can trade on these exchanges without explicitly revealing their identity.
@@
However, in recent years, hybrid models of decentralized exchanges where withdrawals and deposits are permissionless, but certain exchange operations are centralized, have become popular. The most popular exchange of this form is Hyperliquid, which has helped increase the market share of perpetuals volume on DEXs from roughly 1% in 2024 to over 8% in 2025. These hybrid models allow exchanges to use cryptographic commitments to enforce collateral rights while allowing oracles, markets, and solvency to be controlled by a smaller set of participants. This allows the exchange to avoid issues faced by fully open decentralized exchanges, where weak collateral can be added and cannot be sold or liquidated successfully during a market crash to ensure solvency.
More concretely, there have been recent incidents on perpetual DEXs that highlight how novel market structures can be exploited or stressed in practice. These include but are not limited to: Hyperliquid’s XPL pre-market price spike and the JellyJelly attack [46, 90, 89]. These episodes led to emergency safeguards and community post-mortems, underscoring the need for robust oracle design, listing policies, and circuit breakers to mitigate manipulation. Such emergency responses would have been nearly impossible in the fully decentralized setting, demonstrating the value of hybrid models of decentralization in perpetual futures.
@@
Despite the empirical prominence of perpetuals markets, there has been little formal study of the stability and robustness of these markets relative to the large body of research into spot cryptocurrency trading. For instance, the main price stability mechanism used for perpetuals is the funding rate, a continuous payment stream between the long and short positions. There has been some study of how to replicate funding rates via other financial instruments [9, 67, 3], but little to no empirical study of how well such approximations work. Even more surprising is the lack of formal study of how robust perpetual exchanges are under high leverage conditions, despite their usage being driven by high leverage traders.
A crucial component to ensuring the stability of perpetual futures markets with high leverage is liquidation. This is the process of closing a user’s position that is worth less than the cash collateral they have posted to the exchange. Perpetuals exchanges rely on liquidation mechanisms to ensure that they stay solvent — that is, the assets they hold are greater than the liabilities they owe to traders. To the author’s knowledge, there has been no formal study of liquidations in perpetuals exchanges, despite there being a large body of work studying liquidations in overcollateralized cryptocurrency lending (e.g. [93, 75, 111, 98]).
@@
Autodeleveraging is unique in that it algorithmically socializes losses on winners based on a public rule or mechanism that an exchange posts. This means that all of the position closures and implicit liquidations of insolvent positions are performed in an irreversible and atomic manner. The closest analogue to autodeleveraging in traditional finance is a central counterparty (CCP) such as a derivatives clearing house. In traditional finance, CCPs employ default waterfalls and loss mutualization (including variation-margin gains haircutting (VMGH)) that share the same core idea of allocating residual losses across non-defaulting members [49, 96, 45, 44, 77, 116]. CCPs generally have manual, non-algorithmic socialization and disbursement of insurance funds during such events with disputes reconciled via the legal system. We note, furthermore, that only clearing members bare the losses in CCPs, which leads to a different principal-agent dynamic than perpetuals markets as smaller traders need not bare the final losses.
Autodeleveraging, to the best of the author’s knowledge, was first introduced by the crypto exchange Huobi in 2015 [108]. Prior to the implementation of ADL, exchanges would manually socialize losses, with the exchange serving as the sole CCP. By 2016, the BitMEX exchange implemented ADL via a formula that ranked trader positions for ADL by their profit times their leverage [25]. This formula persists to today on virtually every centralized exchange (including Binance, who implemented the formula in 2019 [21]) and on the decentralized exchange Hyperliquid [69]. We note that there are some decentralized exchanges such as Drift [47] and Paradex [91] that use a different ADL mechanism (the pro-rata mechanism defined in §2.4). However, well over 95% of perpetuals trading volume is on exchanges that utilize the BitMEX ranking model.
@@
Inherently, ADL creates moral hazard — that is, losing traders have an incentive to take on more risk in anticipation of being socialized by winning traders. The goal of an exchange is to balance the risk of insolvency against haircutting winning trader profits. Moral hazard is measured by this trade-off — if it is possible to reduce the risk of insolvency to zero without haircutting winning traders, then the system has no moral hazard. On the other hand, if it is not possible to reduce the risk of insolvency by any appreciable amount without severely haircutting winning trades, the system has excessive moral hazard. The job of an exchange operator is to choose an ADL mechanism that tries to induce a low level of moral hazard under most tail event scenarios. This connects to principal–agent models of moral hazard and linear/robust contracting [68, 32, 50], with key differences: many agents act simultaneously, externalities operate through a common solvency constraint, and tail-risk mitigation substitutes for standard effort observability.
@@
Despite the rarity of ADL, it is certainly not hypothetical: there have been dozens of instances of ADL activations on major venues since 2018 [21, 25, 4, 39]. The largest episode of ADL usage was on October 10-11, 2025 (UTC), when outsized liquidations exhausted insurance funds and multiple venues simultaneously invoked ADL [99]. However, other famous incidents such as the March 2020 “Black Thursday” and May 19, 2021 deleveraging event were other prominent days involving ADL usage. Most of these usages of ADL involved controversies, in that highly profitable but lower risk (e.g. lower leverage) traders found their positions closed out by ADL mechanisms. These traders have often threatened to sue exchanges for unfair treatment, with rumors and public speculation surfacing during Oct. 10–11, 2025 (subsequently denied by Wintermute) [41, 114] [^1] See also prior complaints during May 2020 on Binance futures when profitable shorts were closed via ADL [40].
@@
These incidents demonstrate a natural trade-off that perpetuals exchanges have to make: they can either aggressively socialize losses to their winners and (potentially) lose the future revenue of these users while preserving solvency or hold the losses due to insolvency on their balance sheet. In the former case, the exchange effectively chooses to prioritize short-term solvency and potential liquidation profits over long-term value and fees generated by users. The latter case effectively creates a long-term risk for the exchange’s balance sheet, especially if it does not have sufficient funds to cover insolvencies. This strongly suggests that one can view the choice of how an exchange implements ADL as a trade-off between solvency, moral hazard, and long-term revenue for an exchange.
@@
We note that open interest after October 10, 2025 suggests that Hyperliquid has lost nearly 50% of open interest, whereas competitors such as Lighter and even Binance have recovered to pre-event levels. [^2] Source: DeFiLlama perpetuals dashboard, https://defillama.com/perps, accessed Nov. 30, 2025. Open interest is generally more expensive to manipulate than trading volume and can serve as a coarse measurement of expected future revenue for an exchange. Numerous parties have argued that this open interest loss for Hyperliquid is due to their overly aggressive ADL policy. [^3] See, e.g. , public commentary from @fiddybps1, @0xReaz, @Eugene_Bulltime, and @0xLouisT on X (Nov. 2025): https://x.com/fiddybps1/status/1978750722901762321, https://x.com/0xReaz/status/1986486213101166599, https://x.com/Eugene_Bulltime/status/1994380900067582182, https://x.com/0xLouisT/status/1990815109787058654.
@@
    Empirical: Using Hyperliquid’s data from October 10, 2025, where ADL was used over 40 times in a 10-minute period, we evaluate the performance of different ADL mechanisms
@@
Our negative results focus on quantifying limits to how much ADL can actually reduce insolvency without inuring traders with disproportionate losses. We explicitly ask the question of how the loss of profit due to ADL to the trader with the highest PNL scales as the number of positions $n$ grows. We quantify this by looking at ratios of the maximum winner’s profit to the total shortfall (over all traders) and to the worst trader’s shortfall. These ratios can be viewed as analogues of the Value-at-Risk and Expected Shortfall metrics in traditional risk modeling, as we show in Appendix C. The main difference is that while Value-at-Risk and Expected Shortfall measure the total expected loss (over all traders) in tail events, our metrics only measure the expected loss to the most winning trader. Moreover, Appendix C.5 shows that classical VaR/ES rankings echo the same conclusion: queue-based ADL always delivers the lowest top-winner survival even under these traditional risk lenses.
@@
As the negative results show that no ADL mechanism can scale to preserve winner’s profit, a natural follow-up question is, what is the most fair way of socializing profitable users? We consider two notions of fairness in §6: axiomatic and optimization-based. They are both common in the algorithmic fairness literature within machine learning [76, 17, 85, 51, 5, 64]. We demonstrate that the pro-rata ADL rule (used by Drift [47] and Paradex [91]) is the unique fair ADL rule in both the axiomatic and optimization senses [74].
Axiomatic fairness involves defining rules or invariants that a particular ADL mechanism must satisfy to be deemed fair. The rules we analyze for fairness are monotonicity, scale invariance and Sybil resistance. Monotonicity (stable ordering) requires that larger pre-ADL haircutable endowment remains larger post-ADL. Scale invariance states that if everyone scales positions and deficits by an equal factor, relative ADL losses do not change. Sybil resistance (split invariance) states that splitting one economic position across accounts does not change its aggregate ADL seizure.
We show that under mild smoothness conditions, the pro-rata ADL rule is the unique rule to satisfy these axioms. In stark contrast, queue-based mechanisms can violate monotonicity: the greedy waterfall allocation can fully wipe out the top-ranked winner while leaving a nearly identical winner untouched (Appendix D, Proposition D.3). Queue Sybil resistance depends on how accounts are ranked (Appendix D, Proposition D.1): absolute scores (e.g. total PNL) are Sybil vulnerable, while ratio scores such as Hyperliquid’s documented index are Sybil resistant when proportional splitting does not create score ties [72, 1]. In our model and data, the queue-based mechanisms used by Binance and Hyperliquid produce extremely low Profit-to-Solvency Ratios (PTSR), imposing disproportionately large losses on the top winner. This finding contradicts the common intuition that queues are orderly; instead, they concentrate losses on the most successful traders.
@@
A direct extension of the robustness threat model is to consider a sequential ADL problem, where for multiple rounds, a price shock is received, the exchange chooses an ADL policy, and then losses are realized. This model might at first glance seem to only have theoretical value since ADL is itself a rare event. However, on October 10, 2025 there were multiple ADL events that occurred over a short time window on both Hyperliquid and Binance. While the account-level data for Binance’s usage of ADL is not public, we do have this data for Hyperliquid. Starting at 10:21:00 UTC, there were at least 40 recorded ADL events on Hyperliquid in a 10-minute period. This strongly suggests that there is clustering or cascading of ADL events, making the multiple round model necessary to study.
@@
We first show that viewing ADL as a dynamic, multiple round Stackelberg game leads to sharply different equilibria. We prove a crucial “principal-agent” tension: any strategy an exchange executes that minimizes the time to recover solvency (i.e., number of rounds) necessarily maximizes the loss of exchange long-term revenue. This trade-off is the economic heart of the paper: the exchange wants to be solvent in few rounds, but doing so destroys the best customers (e.g. the whales). Intuitively, this occurs because any strategy that quickly recovers solvency autodeleverages the largest winners more aggressively. These winners, however, are also often the most profitable users and generate high long-term fees for the exchange. By liquidating them aggressively, you’re removing the capital that they would use to later create positions and generate fees. We note that threats to sue exchanges for ADL losses in profit after October 10, 2025 provide indirect empirical evidence to support this theoretical result [41, 114].
@@
This implies that any optimal ADL policy over multiple rounds also has to include incentive compatibility constraints that balance future exchange revenue, current exchange solvency, and current trader profits. We construct an explicit algorithm for optimizing a revenue and solvency based objective and show that a simple mirror descent algorithm achieves vanishing regret for this objective. We conclude with a simple example demonstrating that no static policy that dominate the dynamic policy, especially when exchange long-term revenue is a consideration.
Empirical Results: Overshoot Diagnostics.
We conclude by analyzing the October 10, 2025 ADL episode on Hyperliquid under a public-data, partial-observability reconstruction (§9). This event involved roughly \$2.1B of liquidations across a 12-minute cascade. Using loser-side liquidation equity fields (for deficits) and a two-pass no-ADL counterfactual (for realized wealth removed from winners), we estimate an aggregate loser deficit of approximately \$100.1M and a production overshoot vs needed of approximately \$45.0M in PNL dollars at $\Delta=0$, with a short-horizon band of approximately \$45.0M–\$51.7M. Under the same data, a stylized wealth-space queue diagnostic corresponds to approximately \$653.6M in equity dollars. The ratio between these figures reflects winner overcollateralization (equity/PNL $\approx 6.66\times$ trimmed mean across winners).  [^4] Hyperliquid’s on-chain fill log for October 10, 2025 records 34,983 individual ADL executions across 19,337 distinct wallets and 162 tickers. We analyze the per-fill data but aggregate those executions into global time waves (gap 5s) when reporting per-wave totals. Finally, we compare production to transparent benchmark allocations that target the per-wave needed budget $B_{t}^{\mathrm{needed}}$; these benchmarks can match $B_{t}^{\mathrm{needed}}$ closely (subject to capacity caps and contract discreteness), making production overshoot vs needed empirically salient.
Corrections.
In a previous version of this paper, we defined ADL in terms of equity haircuts. After extensive public discussion and critique of how production ADL is executed in contract space (and how that maps to wealth-space impacts), it became clear that haircuts should apply to PNL (profits) rather than total equity (i.e., principal/collateral is protected and senior to the haircutable PNL endowment) [115]. This version reflects these updates: we rewrite the ADL setup with an explicit haircut numéraire (protected cash plus a haircutable endowment, specialized to positive PNL), clarify the observation model used for empirical reconstruction, and show that the theoretical orderings are effectively unchanged while empirical conclusions remain directionally the same with revised magnitudes. This formulation matches the theoretical haircutable endowment definition in §2.